### v2

# Nagaoka ferromagnetism in semiconductor artificial graphene

Abstract
We present the emergence of Nagaoka ferromagnetism in semiconductor-based artificial graphene with realistic Coulomb interaction using high-precision variational and diffusion Monte Carlo methods, complemented by exact diagonalization calculations of the extended Hubbard model. Specifically, we analyze a model of an armchair hexagonal geometry comprising $42$ lattice sites, nanopatterned on GaAs quantum wells with nearest-neighbor distance of $a=50$ nm. Our results reveal a distinct magnetic phase transition driven by the absence/addition of a single electron at half-filling where the ferromagnetic phase is further stabilized by Coulomb scattering terms. We determine effective Hubbard model parameters and identify the magnetic phase transition near $U/t\approx 60$.
Nagaoka ferromagnetism, artificial graphene, graphene quantum dots, quantum simulators, variational Monte Carlo, diffusion Monte Carlo, exact diagonalization
Nagaoka ferromagnetism, originally predicted within the framework of the strongly correlated Hubbard model [1, 2], arises in the infinite U limit when a single hole in a half-filled lattice induces a fully spin-polarized state. Intuitively, the hole motion is fully coherent in the background of a fully polarized spin configuration, minimizing the kinetic energy, in contrast to unpolarized spins that lead to many incoherent paths. Some attempts to generalize Nagaoka’s result lead to a generalized version of Nagaoka’s theorem by Tasaki or the flat-band ferromagnetism of Mielke [3, 4, 5, 6]. Although mathematically rigorous, the Nagaoka theorem and its extensions have limited practical utility in conventional materials, where disorder and competing interactions complicate the picture. To date, suitable conditions for the observation of Nagaoka ferromagnetism have been achieved in a relatively small system—a $2\times 2$ quantum dot plaquette formed electrostatically [7], while Nagaoka polarons were observed in a triangular optical lattice [8]. In previous works, Nagaoka ferromagnetism was investigated within Hubbard model for a single hole for finite and infinite-U limit, and with different hole densities [9, 10, 11, 12, 13, 14, 15], extended Hubbard model [16, 17] and ab initio exact diagonalization [18] for small size systems. Whether Nagaoka ferromagnetism can be realized in larger scale solid-state systems remains an open question as Coulomb interactions include not only finite on-site repulsion but also long-range and scattering terms [19].
Concurrently, solid-state quantum simulators have become an important tool for investigating strongly correlated electron systems, offering unprecedented control over lattice geometries, interactions, and external fields, capabilities that remain challenging to achieve in real materials [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Among solid-state quantum simulators, artificial graphene (AG), recently realized in quantum dot arrays using modulation-doped AlGaAs/GaAs quantum wells [32, 33, 34, 35, 26, 27, 36, 37, 38, 39, 40, 41, 42, 43, 44], has emerged as a powerful platform for exploring tunable Dirac physics and interaction-driven phenomena. The Hubbard parameter in semiconductor AG systems is predicted to be of the order of $U/t\sim 100$ [44], enabling the investigation of correlated insulating phases and emergent magnetism. Thus, with its highly tunable electron filling, interaction strength, and band structure, semiconductor AG provides an ideal platform for probing Nagaoka ferromagnetism under well-controlled conditions, potentially enabling the first direct realization of itinerant ferromagnetism in a programmable quantum system.
In this work, we predict the emergence of Nagaoka ferromagnetism in semiconductor AG, using advanced computational techniques. We show that Nagaoka ferromagnetism can be expected in systems with realistic Coulomb interactions, going beyond Hubbard model. Our findings not only deepen the fundamental understanding of strongly correlated electrons in synthetic lattices but also establish a framework for engineering novel quantum magnetic phases in artificial quantum materials. We employ continuum variational Monte Carlo (VMC) and diffusion Monte Carlo (DMC) methods within the fixed-node approximation to provide a non-perturbative and highly accurate treatment of many-body correlations [45, 46, 47, 48]. We focus on a hexagonal armchair geometry, which serves as an intermediary between finite-size samples and bulk graphene [44]. The system is modeled with a nearest-neighbor distance of $a=50$ nm, in alignment with recent experimental findings [27] and computational studies [44]. Our results reveal a transition from an antiferromagnetic (AFM) to a ferromagnetic (FM) phase which occurs exactly one hole/electron away from the half-filling as described in Nagaoka theorem [1]. Notably, this phase transition appears sharper when considering only the extremal values of $S_{z}$. However, a comprehensive scan over all $S_{z}$ values for a given potential depth and radius demonstrates a more gradual transition.
Our predictions are supported by appropriate exact diagonalization (ED) calculations of the Hubbard model on a hexagonal armchair lattice. Given the prohibitive computational complexity associated with the full Hilbert space for one hole added to the half-filling, the ED calculations are restricted to the subspace corresponding to a single spin flip from a fully spin polarized state. The resulting energy gains from this approach further support the phase transition identified through quantum Monte Carlo (QMC) simulations.
Our model of nanostructured semiconductor AG consists of $N_{e}$ interacting electrons in a honeycomb array of $N$ confining potentials, described by the many-body Hamiltonian in effective atomic units (electronic charge $e$, dielectric constant $\epsilon$, effective mass $m^{*}$, and $\hbar$ are set to 1),
$H=-\frac{1}{2}\sum_{i}^{N_{e}}\nabla^{2}_{i}+\sum_{i}^{N_{e}}V(\textbf{r}_{i})+\sum_{i}^{N_{e}}k|\textbf{r}_{i}|^{2}+\sum_{i<j}^{N_{e}}\frac{1}{r_{ij}}$ (1)
where $1/r_{ij}$ is the Coulomb interaction between the electrons, $V(\textbf{r}_{i})$ is the total potential energy of the confining potentials, and $k$ is the spring constant of quadratic gate potential located at the center of the system which controls the finite size effects. Typical material properties for GaAs, effective electron mass $m^{*}=0.067m_{0}$ and dielectric constant $\epsilon=12.4$, are used. Corresponding effective Bohr radius is $a^{*}_{0}=9.794$ nm, and the effective Hartree energy is $11.857$ meV. The honeycomb array of potential wells is modeled using gaussian-like functions [42, 44],
$V(\textbf{r})=V_{0}\sum\limits_{\textbf{R}_{0}}\exp[-(|\textbf{r}-\textbf{R}_{0}|^{2}/\rho^{2})^{s}]$ (2)
where $V_{0}$ is the potential depth, $\rho$ is the radius and $s$ is the sharpness of the potential wells. $\textbf{R}_{0}$ is the location of the potential wells. In our numerical calculations, dot-to-dot distance was fixed to $a=50$ nm. The sharpness value $s$ was fixed, as well, to $s=1.4$ which is the intermediate point between gaussian ($s=1$) and muffin-tin like ($s=2.8$) potentials [44]. The spring constant of the quadratic gate potential was kept at $k=3.56\times 10^{-4}$ meV/nm2 to reduce finite size effects [44].
Our VMC and DMC calculations were performed using Slater-Jastrow trial wave functions constructed from three different types of orbitals: (i) Gaussian functions on the sites, describing localized states, (ii) Tight-binding (TB) orbitals, suitable for metallic phases with delocalized electrons, and (iii) Mean-field Hubbard (MFH) orbitals, capable of describing both localized and liquid-like states depending on $U/t$. The orbital type yielding to the best fixed-node DMC energy for a given set of system parameters was chosen for further numerical analysis. The details of our VMC and DMC approach can be found in Ref. 44.
Figure 1: Extrapolated spin-spin correlation function plotted against potential depth $V_{0}$ obtained using pair densities for potential radius $\rho=25$ nm. Inset figures are extrapolated spin pair densities in the ground states. (a) $N_{e}=41$, $V_{0}\approx-45$ meV, $S_{z}=41/2$. (b) $N_{e}=42$, $V_{0}\approx-45$ meV, $S_{z}=0$. (c) $N_{e}=41$, $V_{0}\approx-8.84$ meV, $S_{z}=1/2$. (d) $N_{e}=42$, $V_{0}\approx-8.84$ meV, $S_{z}=0$.
For a deep understanding of transition dynamics including magnetic and metallic phases as a function of potential well depth (which controls $U/t$) and electron number, we consider a spin-spin correlation function averaged over all nearest neighbor pairs $(i,j)$, defined as $g=\langle m_{i}m_{j}\rangle/\langle n_{i}n_{j}\rangle$, where $m_{i}$ and $n_{i}$ are the average magnetization and electron density on site $i$ within a radius $r=a/2$. To reveal the internal spin structure, $m_{i}$ and $n_{i}$ are calculated using the pair densities $p_{\sigma\sigma_{0}}(\textbf{r},\textbf{r}_{0})$, the probability of finding an electron with spin $\sigma$ at location r when an electron with spin $\sigma_{0}$ is fixed at location $\textbf{r}_{0}$. The output values of the spin-spin correlation function remain in the $[-1,1]$ range, where $g=-1$ corresponds to an AFM phase, $g=0$ signifies a metallic configuration, and $g=1$ indicates a fully FM phase. In Fig. 1, spin-spin correlation function $g$ is shown as a function of potential depth $V_{0}$ for $N=42$ sites, $\rho=25$ nm and $k\neq 0$. The location of the fixed up-spin electron is denoted by a cross positioned atop a lattice site-strategically chosen to break system symmetry while avoiding edge effects. Both $N_{e}=42$ and $N_{e}=41$ electron numbers are considered for half-filling and a single hole away from half-filling, respectively (see supplemental materials (SM) for QMC $114$ sites results [49]). For these calculations, only maximum and minimum values of $S_{z}$ are considered to determine the ground state of the system. As shown in Fig. 1, the system at the half-filling exhibits a transition from metallic to AFM insulating phase around $V_{0}\approx-18$ meV, and the system with $N_{e}=41$ electrons shows a transition from metallic to FM phase around $V_{0}\approx-30$ meV. The results suggest that the system should make a transition to AFM phase before subtracting a single electron. If the same figure is analyzed vertically, the system exhibits a transition from AFM to FM phase around $V_{0}\approx-30$ meV when an electron is pulled away from the system as it is suggested in Nagaoka theorem. After $V_{0}\approx-30$ meV, the transition maintains its stability. Although, at $\rho=25$ nm, AG exhibits metallic-like behavior as predicted in [44], increasing effect of potential depth $V_{0}$ opens a way to have a localized behavior for the systems. To visualize the phases depicted above by $g$, Figs. 1a-b-c-d show the pair spin densities $p_{\uparrow\uparrow}(\textbf{r},\textbf{r}_{0})-p_{\downarrow\uparrow}(\textbf{r},\textbf{r}_{0})$. In Fig. 1(a), the system have a fully FM phase for $N_{e}=41$ electrons at $V_{0}=-45$ meV with maximum $S_{z}=41/2$ while the system with $N_{e}=42$ electrons at the same $V_{0}$ value Fig. 1(b) exhibits an AFM phase at minimum $S_{z}=0$. Additionally, at the lower extrema value of $V_{0}\approx-8.84$ meV, Fig. 1(c) and 1(d) shows metallic phases for $N_{e}=41$ with $S_{z}=1/2$ and $N_{e}=42$ with $S_{z}=0$, respectively.
Figure 2: (a) Extrapolated spin-spin correlation function plotted against potential depth $V_{0}$ obtained using pair densities for potential radius $\rho=17.5$ nm. (b) Ground state spin $S_{z}$ values plotted against potential depth $V_{0}$. (c) DMC energies plotted against $S_{z}$ at $V_{0}\approx-8.84$ meV for $N_{e}=41$. (d) DMC energies plotted against $S_{z}$ at $V_{0}\approx-27$ meV for $N_{e}=41$. For (a) and (b), the calculations are for $S_{z}$-min/max competition except ”(scan)” results where we consider all $S_{z}$ values.
In Fig. 2, we turn our attention to more localized systems in which potential radius $\rho=17.5$ nm, in order to strengthen the electron correlations [44]. Fig. 2(a) shows spin-spin correlation function as a function of $V_{0}$ for $N_{e}=41$, $N_{e}=42$ and $N_{e}=43$ electrons. In order to validate the results in more detail, a full spin $S_{z}$ scan have been also performed for $N_{e}=41$ to determine the ground states of the system for a given $V_{0}$. In contrast to Fig. 1, half-filling system already shows an AFM behavior starting from higher $V_{0}$ values shown in Fig. 2(a), which is directly correlated with potential radius $\rho=17.5$ nm used in these calculations. The ground states are determined from extrema $S_{z}$ values which we denote as $S_{z}$-min/max. From $S_{z}$-min/max calculations of $N_{e}=41$ electrons, it is observed that the sharp transition from metallic to FM phase is at higher value $V_{0}\approx-15$ meV than $\rho=25$ nm case. Thus, an AFM to FM phase transition also occurs in $\rho=17.5$ nm system depending on electron number yet in a higher $V_{0}$ value effected by strong electron-electron interactions. Full $S_{z}$-scan of $N_{e}=41$ electrons shows a smoother transition (apart from DMC statistical fluctuations) starting from $V_{0}\approx-15$ meV as shown in Fig. 2(a)-(b). The system is in fully polarized FM phase around $V_{0}\approx-26$ meV. In Fig. 2(a), we also show $N_{e}=43$ electrons results in which the system first exhibits a low AFM phase then exhibits a transition to FM state around $V_{0}\approx-23$ meV. While $g$ is slightly lower than $1$ due to the additional opposite spin electron in the system, $S_{z}$ shown in Fig. 2(b) indicates a fully polarized FM phase. Therefore, the AFM to FM phase transition also occurs by adding one electron to the system, as suggested by Nagaoka theorem for an ideal honeycomb system where electron-hole symmetry is present. It is interesting to notice that before a transition to a spin polarized phase after adding one electron, a stable antiferromagnetic order is present that is absent for a one hole case. The asymmetry between $\pm$ one extra electron is expected in a frustrated triangular lattice, and here is related to violation of bipartiteness of the lattice in a realistic system. Weak metallic antiferromagnetism of kinetic origin after adding one hole to a half-filled triangular lattice was predicted by Haerter and Shastry [50].
Fig. 2(c)-(d) show DMC total energies at two examples of extrema $V_{0}$ values of $-8.84$ meV and $-27$ meV for $N_{e}=41$ electrons. MFH $U\approx 20t$ trial wave functions represents more localized states, and TB trial wave functions represents liquid-like states. As seen from Fig. 2(c), the lowest energy is given by the TB trial wave function at $S_{z}=1/2$ for $V_{0}\approx-8.84$ meV. In Fig. 2(d), the ground state is represented by the MFH $U\approx 20t$ trial wave function at $S_{z}=41/2$ for $V_{0}\approx-27$ meV.
Figure 3: Extrapolated spin pair densities for $N_{e}=41$ electrons at several potential depth $V_{0}$ values. (a) $V_{0}\approx-8.84$ meV. (b) $V_{0}\approx-14.89$ meV. (c) $V_{0}\approx-19.74$ meV. (d) $V_{0}\approx-27$ meV.
In Fig. 3, extrapolated spin pair densities for $N_{e}=41$ electrons on $N=42$ sites are shown at several potential depth $V_{0}$ values being chosen to reveal the nature of the phase transition. While a liquid-like state is seen from Fig. 3a at a high $V_{0}$ value, a fully polarized FM state is observed around $V_{0}\approx-27$ meV (Fig. 3d), consistent with $g$ values shown in Fig. 2a. An important point is that before going into a phase transition to FM state, the system exhibits an AFM state around $V_{0}\approx-14.89$ meV due to increasing electron-electron correlations, as seen from Fig. 3b. During the transition around $V_{0}\approx-19.74$ meV in Fig. 3c, the polarization of spins located more around the center. This behavior hints a possible Nagaoka polaron formation in AG systems.
Figure 4: The ratio of onsite Coulomb interaction $U$ to nearest-neighbor hopping parameter $t$, $U/t$, plotted against $V_{0}$ for $\rho=17.5$ nm and $k=3.56\times 10^{-4}$ meV/nm2. (a) $U$ vs. $V_{0}$. (b) $t$ vs. $V_{0}$.
In order to reveal the intrinsic strength of electron-electron interactions, the onsite Coulomb interaction, $U$, is determined using extrapolated electron densities from QMC calculations of $S_{z}$-scanned systems of $N_{e}=41$ electrons. Conversely, the nearest-neighbor hopping parameter, $t$, is estimated through a hybrid approach that combines solutions of the single-particle Schrödinger equation (SPSE) with the diagonalization of the TB Hamiltonian for an artificial benzene molecule in which $\rho=17.5$ nm. The energy spectra obtained from SPSE solutions and TB Hamiltonian diagonalization are systematically compared and fitted. The hopping parameter, $t$, is subsequently extracted by iteratively adjusting and refining the TB model to achieve optimal agreement with the computed spectra. As shown in Fig. 4, the ratio $U/t$ increases in an exponential fashion as $V_{0}$ decreases. The exponential increase of $U/t$ is a direct consequence of increasing $U$ and decreasing $t$ for decreasing $V_{0}$, shown in Fig. 4a and 4b. Around the starting point of AFM to FM phase transition, $V_{0}\approx-15$, the ratio $U/t\approx 60$ while around end point, $V_{0}\approx-26$, the ratio $U/t\approx 170$. Obtained values of the ratio $U/t$ are in good agreement with previous results within triangular optical lattice experiments (for $U/t\approx 72$ Nagaoka polaron around a doublon covers around 30 sites) [8] and within density matrix renormalization group calculation on a square lattice [13].
Figure 5: Exact diagonalization results of the Hubbard Hamiltonian in a subspace of a single spin flip on $N=42$ sites with $N_{e}=41$ electrons (one hole). The energy difference $\Delta E$ between $S=39/2$ and $S=41/2$ states’ lowest energy states. Confining potential strength $\omega=0.5t/a^{2}$ where $a$ is the dot-to-dot distance.
QMC results can be supported by ED calculations in a vicinity of fully spin polarization subspace. ED suffers from an exponential growth of many-body Hilbert space with a system size. For $N=42$ lattice size system in the vicinity of the half-filling, this corresponds to $\sim 10^{23}$ basis states for minimum $S_{z}$ subspace, not accessible within current computational resources. However, one can analyze the stability of a fully spin polarized state from maximum $S_{z}$ subspace ($S^{max}_{z}$) with respect to a spin flip, an energy difference between minimum energies with total spin states with $S=S^{max}-1$ and $S=S^{max}$, $\Delta E=E_{\rm min}(S^{max}-1)-E_{\rm min}(S^{max})$. We consider a hexagonal armchair AG flake consisting of $N=42$ sites in the presence of a quadratic confining potential of strength $\omega=0.5t/a^{2}$, where $a$ is a dot-to-dot distance, a model that mimics QMC calculation in a continuum real space. These computations are carried out within subspaces corresponding to $41$ and $40$ down-spin electrons for a fixed total electron number $N_{e}=41$. ED calculations on a lattice allow us to analyze the role of different Coulomb interaction terms that are all present in real space QMC simulation. In Fig. 5, the energy differences between $S=39/2$ and $S=41/2$ states are presented for four different Hamiltonians with respect to $U/t$ ratio, where $U$ is the onsite Coulomb interaction, $U_{1}$ is the nearest neighbor direct Coulomb interaction, long range consists of all neighbor direct Coulomb interactions, and scattering terms includes assisted-hopping $A$, pair-hopping and exchange terms $X$ (see SM for details [49]). $U_{1}$ is calculated from QMC results as $0.24U$, $A$ is taken as $0.2U$, and the exchange term $X=0.05U$ as in real graphene [51] (these values are also within a parameter range estimated for moire superlattices [52]). For all cases, there is dramatic decrease in the energy difference around $U/t\approx 70$, which is consistent with QMC results. The behavior of energy differences as a function of $U/t$ ratio for only $U$, $U+U_{1}$ and $U+$ long range cases are quite similar, but the transition to a fully spin polarized state is not expected within a range of $U/t$ ratio shown in a figure. However, when scattering terms are added, the energy difference changes sign in which the maximum polarized $S=41/2$ state has the lowest energy around $U/t\approx 60$. This feature is more general regardless of a choice of given parameters, see SM [49]. This hints a possible transition to FM phase around similar $U/t$ ratio as in the QMC results which verifies the phase transition predicted by QMC.
In summary, we demonstrated that Nagaoka ferromagnetism can be probed in semiconductor artificial graphene using accurate variational and diffusion Monte Carlo calculations, alongside exact diagonalization of the Hubbard model. Specifically, quantum Monte Carlo calculations on $N=42$ sites near half-filling reveal an antiferromagnetic to ferromagnetic phase transition driven by the absence of a single electron. Additionally, the phase transition can be induced by adding an electron to the half-filling, consistent with Nagaoka’s theorem on bipartite lattices. As $V_{0}$ decreases (i.e. $U/t$ increases), the $N_{e}=41$ electron system evolves from a liquid-like state to a fully ferromagnetic phase, passing through an intermediate antiferromagnetic phase where spin polarization gradually builds near the center of system, suggesting possible Nagaoka polaron formation. We estimate the $U/t$ ratio and determine that the phase transition occurs above $U/t\approx 60$, indicating that strong electron correlation is necessary. Furthermore, exact diagonalization of the Hubbard model, constrained to near-maximal polarization due to computational limitations, confirms the phase transition. Adding assisted-hopping and exchange interaction terms to the Hamiltonian stabilizes the transition that validates the $U/t$ ratio predictions of quantum Monte Carlo calculations. Thus, our findings establish the presence of an antiferromagnetic to ferromagnetic phase transition triggered by a single electron removal/addition in a solid-state system within a realistic Hamiltonian model fully accounting for long-range and exchange interactions.
Acknowledgements.
The numerical calculations reported in this study were partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).

### v3

@@
 Abstract
-We present the emergence of Nagaoka ferromagnetism in semiconductor-based artificial graphene with realistic Coulomb interaction using high-precision variational and diffusion Monte Carlo methods, complemented by exact diagonalization calculations of the extended Hubbard model.
-Specifically, we analyze a model of an armchair hexagonal geometry comprising $42$ lattice sites, nanopatterned on GaAs quantum wells with nearest-neighbor distance of $a=50$ nm.
-Our results reveal a distinct magnetic phase transition driven by the absence/addition of a single electron at half-filling where the ferromagnetic phase is further stabilized by Coulomb scattering terms.
-We determine effective Hubbard model parameters and identify the magnetic phase transition near $U/t\approx 60$.
+We present the emergence of Nagaoka ferromagnetism in semiconductor-based artificial graphene with realistic Coulomb interaction using high-precision variational and diffusion Monte Carlo methods, complemented by exact diagonalization calculations of the generalized Hubbard model.
+We analyze models of armchair hexagonal geometries nanopatterned on GaAs quantum wells.
+Our results reveal a distinct magnetic phase transition driven by the absence/addition of a single electron at half-filling.
+This form of itinerant magnetism predicted rigorously for Hubbard model remained unascertained in large scale realistic systems.
+We demonstrate that Coulomb scattering terms play a crucial role in stabilizing Nagaoka ferromagnetism, enabling the observation of the phase transition for system parameters near $U/t\approx 60$.
 Nagaoka ferromagnetism, artificial graphene, graphene quantum dots, quantum simulators, variational Monte Carlo, diffusion Monte Carlo, exact diagonalization
@@
 Thus, with its highly tunable electron filling, interaction strength, and band structure, semiconductor AG provides an ideal platform for probing Nagaoka ferromagnetism under well-controlled conditions, potentially enabling the first direct realization of itinerant ferromagnetism in a programmable quantum system.
-In this work, we predict the emergence of Nagaoka ferromagnetism in semiconductor AG, using advanced computational techniques.
-We show that Nagaoka ferromagnetism can be expected in systems with realistic Coulomb interactions, going beyond Hubbard model.
-Our findings not only deepen the fundamental understanding of strongly correlated electrons in synthetic lattices but also establish a framework for engineering novel quantum magnetic phases in artificial quantum materials.
-We employ continuum variational Monte Carlo (VMC) and diffusion Monte Carlo (DMC) methods within the fixed-node approximation to provide a non-perturbative and highly accurate treatment of many-body correlations [45, 46, 47, 48].
-We focus on a hexagonal armchair geometry, which serves as an intermediary between finite-size samples and bulk graphene [44].
+In this work, we predict the emergence of Nagaoka ferromagnetism in semiconductor AG, using advanced computational techniques going beyond Hubbard model.
+We examine a 42-site hexagonal lattice — significantly larger than those studied in prior literature on Nagaoka ferromagnetism, and electrons interacting via long range Coulomb potential, using continuum variational Monte Carlo (VMC) and diffusion Monte Carlo (DMC) methods to provide a non-perturbative and highly accurate treatment of many-body correlations [45, 46, 47, 48].
+We focus on a hexagonal armchair geometry, which serves as an intermediary between finite-size samples and bulk graphene, where edge effects are mitigated through a confining potential [44].
 The system is modeled with a nearest-neighbor distance of $a=50$ nm, in alignment with recent experimental findings [27] and computational studies [44].
 Our results reveal a transition from an antiferromagnetic (AFM) to a ferromagnetic (FM) phase which occurs exactly one hole/electron away from the half-filling as described in Nagaoka theorem [1].
-Notably, this phase transition appears sharper when considering only the extremal values of $S_{z}$.
-However, a comprehensive scan over all $S_{z}$ values for a given potential depth and radius demonstrates a more gradual transition.
 Our predictions are supported by appropriate exact diagonalization (ED) calculations of the Hubbard model on a hexagonal armchair lattice.
 Given the prohibitive computational complexity associated with the full Hilbert space for one hole added to the half-filling, the ED calculations are restricted to the subspace corresponding to a single spin flip from a fully spin polarized state.
-The resulting energy gains from this approach further support the phase transition identified through quantum Monte Carlo (QMC) simulations.
+The ED results further suggest that Coulomb interaction scattering terms are necessary to observe Nagaoka ferromagnetism for realistic system parameters.
 Our model of nanostructured semiconductor AG consists of $N_{e}$ interacting electrons in a honeycomb array of $N$ confining potentials, described by the many-body Hamiltonian in effective atomic units (electronic charge $e$, dielectric constant $\epsilon$, effective mass $m^{*}$, and $\hbar$ are set to 1),
@@
 The location of the fixed up-spin electron is denoted by a cross positioned atop a lattice site-strategically chosen to break system symmetry while avoiding edge effects.
-Both $N_{e}=42$ and $N_{e}=41$ electron numbers are considered for half-filling and a single hole away from half-filling, respectively (see supplemental materials (SM) for QMC $114$ sites results [49]).
+Both $N_{e}=42$ and $N_{e}=41$ electron numbers are considered for half-filling and a single hole away from half-filling, respectively (see supplemental materials (SM) for QMC $N=114$ sites results [49]).
 For these calculations, only maximum and minimum values of $S_{z}$ are considered to determine the ground state of the system.
@@
 We estimate the $U/t$ ratio and determine that the phase transition occurs above $U/t\approx 60$, indicating that strong electron correlation is necessary.
-Furthermore, exact diagonalization of the Hubbard model, constrained to near-maximal polarization due to computational limitations, confirms the phase transition.
+Furthermore, exact diagonalization of the generalized Hubbard model, constrained to near-maximal polarization due to computational limitations, confirms the phase transition.
 Adding assisted-hopping and exchange interaction terms to the Hamiltonian stabilizes the transition that validates the $U/t$ ratio predictions of quantum Monte Carlo calculations.
-Thus, our findings establish the presence of an antiferromagnetic to ferromagnetic phase transition triggered by a single electron removal/addition in a solid-state system within a realistic Hamiltonian model fully accounting for long-range and exchange interactions.
+We note that while QMC calculations were performed for GaAs quantum dots parametrization, similar lattices can also be created in moiré superlattice [53].
+Thus, our findings establish the presence of an antiferromagnetic to ferromagnetic phase transition triggered by a single electron removal/addition in a solid-state system within a realistic Hamiltonian model fully accounting for long-range and scattering interactions.
 Acknowledgements.
-The numerical calculations reported in this study were partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).
+The quantum Monte Carlo calculations reported in this study were performed using the CHAMP program [54] and were partially carried out at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).